Suppose a population of insects according to the law of exponential growth/decay. There were 130 insects after the third day of the experiment and 380 insects after the 7th day. Approximately how many insects were in the original population?


1
Expert's answer
2021-11-29T17:38:51-0500

130 insects after 3 days

380 insects after 7 days

"y=Ae^{Bt}"

"130=Ae^{3B}" ..........(i)

"380=Ae^{7B}" ..........(ii)

Dividing equation (ii) by (i) we get;

"\frac{380}{130}=\frac{Ae^{7B}}{Ae^{3B}}"

"\frac{38}{13}=\frac{e^{7B}}{e^{3B}}=e^{7B}*e^{-3B}=e^{4B}"

"ln(\frac{38}{13})={4B}"

B=0.2681

Replace this in equation (i) to get A

"130=Ae^{3B}"

"130=Ae^{3*0.2681}=Ae^{0.80}"

130=2.226A

A=58.4

Initial insect population at t=0

"y=Ae^{B*0}"

"\therefore y=A"

y=58.4 insects

"\approx 59" insects